# TabReD: A Benchmark of Tabular Machine Learning in-the-Wild

### v1

@@ 3.2 Constructing the TabReD Benchmark
In this section, we introduce the new Tabular benchmark with Real-world industrial Datasets (TabReD). To construct TabReD we source datasets from Kaggle competitions and industrial ML applications at a large tech company [^5] Yandex. We adhere to the following criteria when selecting datasets for TabReD. (1) Datasets should be inherently tabular, as discussed in subsection 3.1 and Kohli et al. [34]. (2) Feature engineering and feature collection efforts should be closer to the industry practices. We adapt feature-engineering code by studying competition forums for Kaggle datasets, and we use the exact features from production ML systems for the newly introduced datasets. (3) Exclude datasets with known data leakage, we also take care to avoid leakage in the newly introduced datasets. (4) Datasets should have timestamps available and should have enough samples for the time-based train/test split (this excludes datasets, where future instances are not publicly available, making it impossible to do local validation for research purposes correctly).
The table with our annotations of Kaggle competitions is available in the Appendix C. Below we provide short descriptions of datasets and corresponding tasks. [^6] For detailed information regarding data composition, preprocessing and licenses see supplementary materials. We also provide a brief overview for each dataset in Appendix D.
Homesite Insurance. This is a dataset from a Kaggle competition hosted by Homesite Insurance [14]. The task is predicting whether a customer will buy a home insurance policy based on user and insurance policy features (user, policy, sales and geographic information). Each row in the dataset corresponds to a potential [customer, policy] pair, the target indicates whether a customer bought the policy.
Ecom Offers. This is a dataset from a Kaggle competition hosted by the online book and game retailer DMDave [15]. The task in this dataset is a representative example of modeling customer loyalty in e-commerce. Concretely, the task is classifying whether a customer will redeem a discount offer based on features from two months’ worth of transaction history. We base our feature engineering on one of the top solutions [40].
HomeCredit Default. This is a second iteration of the popular HomeCredit tabular competition [23]. The task is to predict whether bank clients will default on a loan, based on bank internal information and external information like credit bureau and tax registry data. This year competition focus was the model prediction stability over time. Compared to the more popular prior competition, this time there is more data and the timestamps are available. We base feature engineering and preprocessing code on top solutions [31].
Sberbank Housing. This dataset is from a Kaggle competition, hosted by Sberbank [3]. This dataset provides information about over 30000 transactions made in the Moscow housing market. The task is to predict the sale price of each property using the provided features describing each property condition, location, and neighborhood, as well as country economic indicators at the moment of the sale. We base our preprocessing code on discussions and solutions from the competition [4, 1].
3.2.1 New Datasets From In-the-Wild ML Applications
Here, we describe the datasets used by various ML applications that we publish with TabReD. All of these datasets were preprocessed for later use by a model in production ML systems. We apply deterministic transforms to anonymize the data for some datasets. We only publish the preprocessed data, as the feature engineering code and internal logs are proprietary. We provide further details regarding licenses, preprocessing and data composition in the datasheet with supplementary materials.
Maps Routing. For this dataset, the task is to predict the travel time in the car navigation system based on the current road conditions. The features are aggregations of the road graph statistics for the particular route and various road details (like speed limits). The target is the logarithm of seconds per kilometre.
Weather. For this dataset, the task is weather temperature forecasting. This is a dataset similar to the one introduced in [38], except it is larger, and the time-based split is available and used by default. The features are from weather station measurements and weather forecast physical models. The target is the true temperature for the moment in time.
4 Experiments
In this section, we demonstrate the utility of the introduced TabReD benchmark, by answering the following questions:
  • Q1 (subsection 4.1) How do recent tabular DL advances on academic benchmarks transfer to the industry setting facilitated by TabReD?
  • Q2 (subsection 4.2) How does accounting for a temporal shift in the data affects model performance and ranking in tabular machine learning?
@@ 4 Experiments
Experimental setup. We adopt training, evaluation and tuning setup from [21]. We tune hyperparameters for most methods [^7] Trompt is the only exception, due to the method’s time complexity. For Trompt we evaluate the default configuration proposed in the respective paper. using Optuna from Akiba et al. [2], for DL models we use AdamW optimizer and optimize MSE loss or binary cross entropy depending on the datasets. We keep our training protocol simple and do not employ additional training techniques for DL models, like pretraining, custom loss functions, augmentations and many others. We randomly subsample large datasets (Homecredit Default, Cooking Time, Delivery ETA and Weather) to make more extensive hyperparameter tuning feasible. For other details regarding data preprocessing, dataset statistics, tuning protocols and hyperparameters, see Appendix B. Below, we describe a set of models we evaluate on the proposed benchmark.
@@
Non DL Baselines. We include three main implementations of Gradient Boosted Decision Trees: XGBoost [13], LightGBM [29] and CatBoost [41], as well-established non-DL baselines for tabular data prediction. We also include Random Forest [9] and linear model as the basic simple ML baselines.
Baseline DL Models. We include two baselines from [19] – MLP and FT-Transformer. We use MLP as the simplest DL baseline, FT-Transformer as a representative attention-based tabular DL model, which is often claimed to be state-of-the-art [e.g. 19, 46, 36, 10]. In addition, we include DCNv2 as it was repeatedly used in real-world production settings as reported by [52, 5]. We hypothesize that the presence of a feature cross-module could bring performance improvements on the benchmark where datasets have more features. We also test alternative MLP-like backbones in ResNet [19] and SNN [33].
Numerical Features Embeddings. We include MLP with embeddings for numerical features from Gorishniy et al. [20]. It was demonstrated that numerical embeddings provide considerable performance improvements on academic datasets, and make simple MLP models compete with attention-based models. We find this technique simple and effective, thus important to evaluate in a new setting.
@@
Trompt. We include a novel tabular DL model from Chen et al. [12], which was shown to outperform Transformer for tabular data variants [46, 19] on the benchmark from Grinsztajn et al. [22]. Its strong performance on an established academic benchmark aligns well with our goal of finding out how results obtained on academic benchmarks might generalize to TabReD.
TabR. A recent addition to the tabular DL model arsenal from Gorishniy et al. [21], a retrieval-based model, demonstrating impressive performance on commonly used academic benchmarks [19, 22]. We evaluate the TabR-S variation, which does not include numerical embeddings, as we test the efficacy of numerical embeddings via MLP-PLR, mentioned above. Note that, during training, for a given training object, the vanilla TabR can retrieve nearest neighbors from “future” training objects (i.e. from training objects with greater timestamps). In the light of the temporal shifts of the TabReD datasets, this can be an issue, since the retrieval from future will not be possible for test data. Moreover, even without that issue, the similarities between future test objects and past training objects can still be too different from what TabR sees during training. As a small remedy to these issues, we introduce TabR-S (causal): during training, for a given object, this version of TabR retrieves nearest neighbors only “from the past” (from “older” datapoints), which is more aligned with how TabR will make predictions for the test data. We hypothesize that a more complete solution could be to immediately add “older” test objects as nearest neighbor candidates for more recent test objects. This would utilize the unique ability of retrieval-based models to accomodate new data without retraining, which comes at a price of the increased implementation complexity.
OOD Robustness Methods. We also evaluate two methods that aim to mitigate the effect of distribution shift. The first one is DeepCORAL [48], we adapt the method to the temporal shift setting by bucketing timestamps into different domains, similar to Wild-Time [53]. The second method is Deep Feature Reweighting (DFR) [32], we adapt the method by finetuning the representation of the MLP baseline on the latter instances of the train dataset.
4.1 How do recent tabular DL advances transfer to a more realistic evaluation setting?
We evaluate all abovementioned methods to answer Q1. Results are summarized in Table 2.
@@ 4.1 How do recent tabular DL advances transfer to a more realistic evaluation setting?
Table 2: Performance comparison of tabular ML models on new datasets. Bold entries represent the best methods on each dataset, with standard deviations over 15 seeds taken into account. The last column contains algorithm rank averaged over all datasets (for details, see the subsection B.2). The ranks in bold correspond to the top-3 classical ML methods and the top-3 DL methods.
@@
XGBoost 0.9601 0.5763 0.8670 0.2419 0.4823 0.5468 0.1616 1.4671 2.2 $\pm$ 1.3
LightGBM 0.9603 0.5758 0.8664 0.2468 0.4826 0.5468 0.1618 1.4625 2.6 $\pm$ 1.2
CatBoost 0.9606 0.5596 0.8621 0.2482 0.4823 0.5465 0.1619 1.4688 2.8 $\pm$ 1.4
RandomForest 0.9570 0.5764 0.8269 0.2640 0.4884 0.5959 0.1653 1.5838 6.4 $\pm$ 1.7
Linear 0.9290 0.5665 0.8168 0.2509 0.4882 0.5579 0.1709 1.7679 7.4 $\pm$ 2.2

### v4

@@ 4 Constructing the TabReD benchmark
In this section, we introduce the new Tabular benchmark with Real-world industrial Datasets (TabReD). To construct TabReD we utilize datasets from Kaggle competitions and industrial ML applications at a large tech company. We adhere to the following criteria when selecting datasets for TabReD. (1) Datasets should be inherently tabular, as discussed in section 3 and Kohli et al. (2024). (2) Feature engineering and feature collection efforts should be as close as possible to the industry practices. We adapt feature-engineering code by studying competition forums for Kaggle datasets, and we use the exact features from production ML systems for the newly introduced datasets. (3) We also take care to avoid leakage in the newly introduced datasets. (4) Datasets should have timestamps available and should have enough samples for the time-based train/test split.
We summarize the main information about our benchmark in Table 2. The complete description of each included dataset can be found in Appendix B. The analysis of feature collinearity is provided in subsection A.2. We also provide the table with our annotations of Kaggle competition, used to filter datasets from Kaggle Appendix D.
Table 2: Short description of datasets in TabReD. Numbers in parentheses denote full data size. We use random subsets of large datasets to make extensive hyperparameter tuning feasible.
Dataset # Samples # Features Source Task Description
Sberbank Housing 20K 387 Kaggle Real estate price prediction
Homesite Insurance 224K 296 Kaggle Insurance plan acceptance prediction
Ecom Offers 106K 119 Kaggle Predict whether a user will redeem an offers
HomeCredit Default 381K (1.5M) 696 Kaggle Loan default prediction
Cooking Time 228K (10.6M) 195 New Restaurant order cooking time estimation
Delivery ETA 224K (6.9M) 225 New Grocery delivery courier ETA prediction
Maps Routing 192K (8.8M) 1026 New Navigation app ETA from live road-graph features
Weather 605K (6.0M) 98 New Weather prediction (temperature)
4.1 On the role and limitations of TabReD
We see our TabReD benchmark as an important addition to the landscape of tabular datasets. TabReD has the properties that are underrepresented in current benchmarks, including the time-based splits and the “feature-rich” data representation. And, as we will show in section 5, these properties become a non-trivial challenge for some tabular models and techniques. Thus, TabReD enables an additional evaluation in the industrial-like setting that can help in identifying limitations of novel methods. As a bonus point, TabReD increases the number of publicly available real-world datasets with more than one million objects.
However, TabReD is not a full replacement for current benchmarks, and has certain limitations. First, it is biased towards industry-relevant ML applications, with large sample sizes, extensive feature engineering and temporal data drift. Second, TabReD does not cover some important domains such as medicine, science, and social data. Finally, we note that the lack of precise feature information may limit some potential future applications of these datasets, like leveraging feature names and descriptions with LLMs.
5 How do tabular dl techniques transfer to TabReD conditions?
In this section, we use TabReD to analyze how tabular deep learning advances on previously available benchmarks transfer to the temporally evolving, feature-rich industrial setting. In subsection 5.1 we introduce our experimental setup and list techniques and baselines considered in our study. In subsection 5.2 we evaluate all these techniques on TabReD and discuss the results. In subsection 5.3 we contrast results on commonly used datasets with TabReD and show that some techniques useful on previous benchmarks are less effective on TabReD than others. In subsection 5.4 we analyze the influence of temporal shift on evaluation results and method comparison. We also study methods of increasing robustness to distribution shift on TabReD in subsection A.3.
5.1 Experimental setup and tabular deep learning techniques
@@
Experimental setup. We adopt training, evaluation and tuning setup from Gorishniy et al. (2024). We tune hyperparameters for most methods [^5] Trompt is the only exception, due to the method’s time complexity. For Trompt we evaluate the default configuration proposed in the respective paper. using Optuna from Akiba et al. (2019), for DL models we use the AdamW optimizer and optimize MSE loss or binary cross entropy depending on the dataset. By default, each dataset is temporally split into train, validation and test sets. Each model is selected by the performance on the validation set and evaluated on the test set (both for hyperparameter tuning and early-stopping). Test set results are aggregated over 15 random seeds for all methods, and the standard deviations are taken into account to ensure the differences are statistically significant. We randomly subsample large datasets (Homecredit Default, Cooking Time, Delivery ETA and Weather) to make more extensive hyperparameter tuning feasible. For extended description of our experimental setup including data preprocessing, dataset statistics, statistical testing procedures and exact tuning hyperparameter spaces, see Appendix C. Below, we describe the techniques we evaluate on TabReD.
@@
Non DL Baselines. We include three main implementations of Gradient Boosted Decision Trees: XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017) and CatBoost (Prokhorenkova et al., 2018), as well-established non-DL baselines for tabular data prediction. We also include Random Forest (Breiman, 2001) and linear model as the basic simple ML baselines to ensure that datasets are non-trivial and are not saturated by simplest baselines.
Tabular DL Baselines. We include two baselines from (Gorishniy et al., 2021) – MLP and FT-Transformer. We use MLP as the simplest DL baseline and FT-Transformer as a representative baseline for the attention-based tabular DL models. Attention-based models are often considered state-of-the-art (e.g. Gorishniy et al., 2021; Somepalli et al., 2021; Kossen et al., 2021; Chen et al., 2023a) in tabular deep learning research.
@@
Other Tabular DL Models. In addition to baseline methods, we include DCNv2 as it was repeatedly used in real-world production settings as reported by (Wang et al., 2020; Anil et al., 2022). We also test alternative MLP-like backbones in ResNet (Gorishniy et al., 2021) and SNN (Klambauer et al., 2017). We also include Trompt (Chen et al., 2023b), it was shown to outperform Transformer for tabular data variants (Somepalli et al., 2021; Gorishniy et al., 2021) on the benchmark from Grinsztajn et al. (2022). Its strong performance on an established academic benchmark aligns well with our goal of finding out how results obtained on previously available benchmarks generalize to TabReD.
Numerical Feature Embeddings. We include MLP with embeddings for numerical features from Gorishniy et al. (2022). Numerical embeddings provide considerable performance improvements, and make simple MLP models compete with attention-based models and GBDTs. Furthermore, the success of this architectural modification was replicated on benchmarks in Ye et al. (2024) and Holzmüller et al. (2024). We find this simple, effective and proven technique important to evaluate in a new setting.
Retrieval-based models are a recent addition to the tabular DL model arsenal. We evaluate TabR from Gorishniy et al. (2024) and ModernNCA from Ye et al. (2024). Both retrieval-based models demonstrate impressive performance on common datasets from Gorishniy et al. (2021); Grinsztajn et al. (2022) and outperform strong GBDT baselines with a sizeable margin, which warrants their evaluation on TabReD. We exclude numerical embeddings from both models to test the efficacy of retrieval in a new setting in isolation.
Improved Training Methodologies like the use of cut-mix like data augmentations (Somepalli et al., 2021; Chen et al., 2023a) or auxiliary training objectives with augmentations and long training schedules (Rubachev et al., 2022; Lee et al., 2024) have produced considerable gains on current benchmarks. We include two methods from Rubachev et al. (2022), namely long training with data-augmentations “MLP aug.” and a method with an additional reconstruction loss “MLP aug. rec.” [^6] In the original work the methods are called “MLP sup” and “MLP rec-sup”, we use subscript aug to highlight that methods use augmentations and differentiate them from traditional supervised baselines.
Ensembles are considered a go-to solution for improving performance in many ML competitions. Ensembles have also shown effectiveness in tabular DL (Gorishniy et al., 2021; Shwartz-Ziv & Armon, 2021). We test this technique on TabReD as well.
5.2 Results
In this section, we evaluate all techniques outlined above. Results are summarized in Table 3. Below, we highlight our key takeaways.
@@ 5.2 Results
Table 3: Performance comparison of tabular ML models on new datasets. Bold entries represent the best methods on each dataset, with standard deviations over 15 seeds taken into account. The last column contains algorithm rank averaged over all datasets (for details, see the subsection C.2). The ranks in bold correspond to the top-3 classical ML methods and the top-3 DL methods.
@@
XGBoost 0.9601 0.5763 0.8670 0.2419 0.4823 0.5468 0.1616 1.4671 2.9 $\pm$ 1.5
LightGBM 0.9603 0.5758 0.8664 0.2468 0.4826 0.5468 0.1618 1.4625 3.1 $\pm$ 1.5
CatBoost 0.9606 0.5596 0.8621 0.2482 0.4823 0.5465 0.1619 1.4688 3.4 $\pm$ 1.7
RandomForest 0.9570 0.5764 0.8269 0.2640 0.4884 0.5959 0.1653 1.5838 7.8 $\pm$ 2.0
Linear 0.9290 0.5665 0.8168 0.2509 0.4882 0.5579 0.1709 1.7679 8.8 $\pm$ 2.7
Tabular DL Models
MLP 0.9500 0.6015 0.8545 0.2508 0.4820 0.5504 0.1622 1.5470 5.0 $\pm$ 1.8
SNN 0.9492 0.5996 0.8551 0.2858 0.4838 0.5544 0.1651 1.5649 6.6 $\pm$ 2.0
DCNv2 0.9392 0.5955 0.8466 0.2770 0.4842 0.5532 0.1672 1.5782 7.6 $\pm$ 2.3
ResNet 0.9469 0.5998 0.8493 0.2743 0.4825 0.5527 0.1625 1.5021 5.8 $\pm$ 2.0
FT-Transformer 0.9622 0.5775 0.8571 0.2440 0.4820 0.5542 0.1625 1.5104 4.8 $\pm$ 1.6
MLP (PLR) 0.9621 0.5957 0.8568 0.2438 0.4812 0.5527 0.1616 1.5177 3.8 $\pm$ 1.4
Trompt 0.9588 0.5803 0.8355 0.2509 0.4809 0.5519 0.1624 1.5187 5.4 $\pm$ 2.1
Ensembles
MLP ens. 0.9503 0.6019 0.8557 0.2447 0.4815 0.5494 0.1620 1.5186 4.1 $\pm$ 2.0
MLP-PLR ens. 0.9629 0.5981 0.8585 0.2381 0.4806 0.5518 0.1612 1.4953 2.4 $\pm$ 1.5
Training Methodologies
MLP aug. 0.9523 0.6011 0.8449 0.2659 0.4832 0.5532 0.1631 1.5193 6.0 $\pm$ 2.0
MLP aug. rec. 0.9531 0.5960 0.7453 0.2515 0.4834 0.5541 0.1636 1.5160 6.8 $\pm$ 2.8
Retrieval Augmented Tabulard DL
TabR 0.9487 0.5943 0.8501 0.2820 0.4828 0.5514 0.1639 1.4666 6.0 $\pm$ 2.2
ModernNCA 0.9514 0.5765 0.8531 0.2593 0.4825 0.5498 0.1625 1.5062 5.6 $\pm$ 1.6
GBDT and MLP with embeddings (MLP-PLR) are the overall best models on the TabReD benchmark. These findings suggest that numerical feature embeddings (Gorishniy et al., 2022), which have shown success on existing benchmarks, maintain their utility in the new evaluation scenario. Ensembles also bring consistent performance improvements to MLP and MLP-PLR in line with the existing knowledge on prior benchmarks and practice.
FT-Transformer is a runner-up, however, it can be slower to train because of the attention module that causes quadratic scaling of computational complexity w.r.t. the number of features. The latter point is relevant for TabReD, since the TabReD datasets have more features than an average dataset used in literature.
SNN, DCNv2, ResNet and Trompt are no better than the MLP baseline. Although Trompt showed promising results on a benchmark from Grinsztajn et al. (2022), results on TabReD are mixed and worse than MLP on average. Furthermore, efficiency-wise, Trompt is significantly slower than MLP, and even slower than FT-Transformer.
Retrieval-Based Models prove to be less performant on TabReD. One notable exception is the Weather dataset, where TabR has the second-best result. ModernNCA is closer to the MLP baseline on average, but its benifits seen on previously available benchmarks do not transfer to the TabReD setting either. We expand on potential challenges TabReD presents for retrieval-based models in the following section.
Improved Training Methodologies. Better training recipes that leverage long pre-training on target datasets mostly do not transfer to the setting presented by TabReD. Homesite Insurance and Weather datasets are the only exceptions, where training recipes bring some performance improvements.
5.3 Comparison with prior benchmarks
To further illustrate the utility of the TabReD benchmark for differentiating tabular DL techniques, we compare results for a range of recently proposed methods on an existing benchmark from Gorishniy et al. (2024) and on TabReD. We look at four techniques in this section: improved model architectures (MLP-PLR and XGBoost as a reference classic baseline), ensembling, improved training methodologies (“MLP aug.”, “MLP aug. rec.”) and retrieval-based models (ModernNCA, TabR). The results are in Figure 1. We summarize our key observations below.
Figure 1: Comparison of tabular DL algorithmic improvements on TabReD and on a popular benchmark. We plot average relative percentage improvement over the MLP baseline on both benchmarks. Ensembling and Numerical Embeddings successfully transfer to TabReD. However, success of retrieval-based models and improved training methods does not translate to the proposed setting.
We can see that ensembling and embeddings for numerical features are beneficial on both benchmarks. But two remaining techniques: improved training recipes and retrieval-based models underperform on TabReD, while showing consistently high performance on previously available datasets.
We hypothesize that the two TabReD dataset characteristics are at play here. First, feature complexities like multicollinearity and presence of noisy features of more feature-rich TabReD datasets might affect neighbors selection and the feature-shuffle augmentation (for a comparison of feature complexities, see subsection A.2). Second, retrieval-based models rely on an assumption that train objects (used in retrieval) are useful for predicting test instances. This can be violated by the presence of gradual temporal shift. A similar argument may explain the underperformance of long training recipes, as those models might also implicitly “memorize” harder train-set examples through longer training (Rahaman et al., 2019; Feldman & Zhang, 2020). Understanding these phenomena is an interesting avenue for future research.
Summary. Overall, the above results reinforce TabReD as an effective tool for uncovering previously unnoticed failure modes of tabular methods.
5.4 Influence of data validation splits on model ranking
In this section, we investigate the importance of taking temporal shift in tabular datasets into account in evaluation. For this we create three time-based splits (with a sliding window over all samples) and three corresponding randomly shuffled splits, keeping train, validation and test sizes the same. We average all results over 15 random initialization seeds and three data splits. We are interested in methods ranking and relative performance differences between results on random and time-based test splits. For this experiment, we consider MLP, MLP-PLR, XGBoost and TabR-S. This selection covers multiple different paradigms: retrieval-based models, parametric DL and strong non-DL baselines. Furthermore, those methods have diverse performance on our benchmark (see Table 3). Results are summarized in Figure 2. Below, we highlight key takeaways.
Temporal Shift’s Influence on Performance. We see that the spread of scores depending on random initialization for each model and data split is generally larger on the time-split based test set (most notable difference is on Sberbank Housing, Delivery ETA and Ecom Offers). This hints that temporal shift is present in the proposed datasets (see extended discussion in the subsection A.1 on the presence of shifts in TabReD datasets). As the random splits are commonplace in current benchmarks, they could present overly optimistic performance estimates to what one might expect in real-world application scenarios.
Temporal Shift’s Influence on Ranking and Relative Difference. We can see that the ranking of different model categories and the spreads between model performance scores change when we use randomly split test sets. One notable example is XGBoost decreasing its performance margin to MLPs when evaluated on temporally shifted test sets (Sberbank Housing, Cooking Time, Delivery ETA, Weather, Ecom Offers and Quote Conversion datasets – most clearly seen comparing MLP-PLR with XGBoost). This might indicate that GBDTs are less robust to shifts, conversely performing better on random splits, by possibly exploiting time-based leakage. Another notable example is TabR-S outperforming the baseline MLP (Cooking Time and Homesite Insurance) and even XGBoost (Weather). These findings prove that curating datasets that represent real-world use cases is important for the continued stable progress in the field of tabular machine learning and further research adoption.
6 Future Work
We have demonstrated the importance of taking temporal shifts into account and benchmarked a wide range of prominent tabular DL techniques on TabReD. However, there are still many research questions and techniques like continual learning, gradual temporal shift mitigation methods, missing data imputation and feature selection, that could be explored with TabReD. The question of why some effective techniques fail to transfer to TabReD is worth further investigation, we posit that two key data characteristics in temporal data drift and feature complexities like multicollinear or noisy features are at play here.
7 Conclusion
In this work, we analyzed the existing tabular ML datasets typically used in the literature and identified their limitations. Also, we described two conditions common in typical deployment scenarios that are underrepresented in current benchmarks. We then composed a new benchmark TabReD that closely reflects these conditions and follows the best industrial practices. We carefully evaluated many recent tabular DL developments in TabReD settings and found that simple baselines like MLP with embeddings and deep ensembles, as well as GBDT methods such as XGBoost, CatBoost and LightGBM work the best, while more complicated tabular DL methods fail to transfer their increased performance from previous academic benchmarks. We believe that TabReD can serve as an important step towards more representative evaluation and will become a testing ground for future methods of tabular DL.
Reproducibility Statement
We describe our experimental setup in subsection 5.1 and Appendix C. We also provide code with the submission. Dataset downloading and preprocessing is handled by the provided code (see the ./preprocessing folder. Newly introduced datasets would be available upon acceptance and are not available for anonymity. Instructions to reproduce experiments and plots are in the provided code README.md
Appendix A Additional Results and Artifacts
A.1 Exploration of TabReD datasets: shifts
In this section, we explore the temporal shift aspect of the TabReD datasets.
We plot standard deviations of MLP ensemble predictions over time together with model errors over the same timeframe. The plots are in Figure 3. These plots show a more nuanced view of the distribution shift in TabReD, and it’s relationship to model performance:
Figure 3: Relationship between distribution shift and performance on a subset of TabReD datasets. We use std in ensemble of MLP predictions as a proxy for the distribution shift. On the right side, we show errors (MAE for regression and error rate for binary classification)
  • Some datasets exhibit trends where the shift increases over time and the error rate increases with it (this can be seen on Homesite Insurance and Sberbank Housing).
  • In addition to global shift over time, some datasets exhibit strong seasonal behavior (e.g. there are specific times of day in the test set where performance drops significantly on the Cooking Time dataset).
  • However, sometimes data shift might lead to seemingly better performance over time (e.g. as time goes by model predictions might improve). As variance and irreducible noise in the target variable can decrease over time, it could cause the performance metrics to improve over time. It is important to note that the model still suffers from the detrimental effects of the distributional shift, since if it was trained on the examples from the same domain as the one comprising the OOD test set, the performance would have been even better). The evolution of the degree of data shift and model errors on Homecredit Default is an example of such behavior.
A.2 Exploration of TabReD datasets: feature correlations
In this section, we explore the “feature-rich” aspect of TabReD datasets. For this, we plot the linear feature correlations and unary feature importances. To compute feature importances we use the mutual information in sklearn (Pedregosa et al., 2011) using methods from Kraskov et al. (2004).
Correlations together with feature-target mutual information for two benchmarks are in Figure 4.
Figure 4: Feature correlations and importance via mutual information with target. On the left are datasets from Gorishniy et al. (2024), on the right are TabReD datasets. Datasets on the right are clearly more complex in terms of number of features and their correlation and importance patterns. The only comparably complex dataset on the left is Microsoft.
A.3 Distribution shift robustness methods
We evaluate two methods that aim to mitigate the effect of distribution shift. The first one is DeepCORAL (Sun & Saenko, 2016), we adapt the method to the temporal shift setting by bucketing timestamps into different domains, similar to Wild-Time (Yao et al., 2022). The second method is Deep Feature Reweighting (DFR) (Kirichenko et al., 2023), we adapt the method by finetuning the representation of the MLP baseline on the latter instances of the train dataset.
Table 4: Study of distribution shift robustness methods on TabReD.
Methods Classification (ROC AUC ↑) Regression (RMSE ↓) Average Rank
MLP 0.9500 0.6015 0.8545 0.2508 0.4820 0.5504 0.1622 1.5470 1.0 $\pm$ 0.0
CORAL 0.9498 0.6004 0.8549 0.2645 0.4821 0.5498 0.1622 1.5591 1.4 $\pm$ 0.7
DFR 0.9499 0.6013 0.8545 0.2494 0.4819 0.5515 0.1626 1.5513 1.4 $\pm$ 0.5
Both DFR and DeepCORAL do not improve upon the MLP baseline, in line with recent work by Gardner et al. (2023); Kolesnikov (2023) for other distribution shifts.
Appendix B TabReD dataset details
In this section, we provide more detailed dataset descriptions.
B.1 Datasets sourced from Kaggle competitions
Below, we provide short descriptions of datasets and corresponding tasks.
Homesite Insurance. This is a dataset from a Kaggle competition hosted by Homesite Insurance (Darrel, 2015). The task is predicting whether a customer will buy a home insurance policy based on user and insurance policy features (user, policy, sales and geographic information). Each row in the dataset corresponds to a potential [customer, policy] pair, the target indicates whether a customer bought the policy.
Ecom Offers. This is a dataset from a Kaggle competition hosted by the online book and game retailer DMDave (DMDave, 2014). The task in this dataset is a representative example of modeling customer loyalty in e-commerce. Concretely, the task is classifying whether a customer will redeem a discount offer based on features from two months’ worth of transaction history. We base our feature engineering on one of the top solutions (MLWave, 2014).
HomeCredit Default. This is a second iteration of the popular HomeCredit tabular competition (Herman et al., 2024). The task is to predict whether bank clients will default on a loan, based on bank internal information and external information like credit bureau and tax registry data. This year competition focus was the model prediction stability over time. Compared to the more popular prior competition, this time there is more data and the timestamps are available. We base feature engineering and preprocessing code on top solutions (Kim, 2024).
Sberbank Housing. This dataset is from a Kaggle competition, hosted by Sberbank (Alexey Matveev, 2017). This dataset provides information about over 30000 transactions made in the Moscow housing market. The task is to predict the sale price of each property using the provided features describing each property condition, location, and neighborhood, as well as country economic indicators at the moment of the sale. We base our preprocessing code on discussions and solutions from the competition (Alijs & Johnpateha, 2017; Sidorova, 2017).
B.2 New datasets from in-the-wild ml applications
Here, we describe the datasets used by various ML applications that we publish with TabReD. All of these datasets were preprocessed for later use by a model in production ML systems. We apply deterministic transforms to anonymize the data for some datasets. We only publish the preprocessed data, as the feature engineering code and internal logs are proprietary. We will provide further details regarding licenses, preprocessing and data composition in the datasheet with supplementary materials upon acceptance (we can’t disclose dataset details due to submission anonymity).
Maps Routing. For this dataset, the task is to predict the travel time in the car navigation system based on the current road conditions. The features are aggregations of the road graph statistics for the particular route and various road details (like speed limits). The target is the logarithm of seconds per kilometer.
Weather. For this dataset, the task is weather temperature forecasting. This is a dataset similar to the one introduced in (Malinin et al., 2021), except it is larger, and the time-based split is available and used by default. The features are from weather station measurements and weather forecast physical models. The target is the true temperature for the moment in time.
Appendix C Extended experimental setup description
This section is for extended information on the experimental setup.
C.1 Source code
We include source code for reproducing the results in the supplementary material archive with a brief README.md with instructions on reproducing the experiments. We also publish code for dataset preparation.
C.2 Tuning, evaluation, and model comparisons
In tuning and evaluation setup, we closely follow the procedure described in (Gorishniy et al., 2021; 2024).
When comparing the models, we take standard deviations over 15 random initializations into account. We provide the ranks each method achieved below. In Table 3, following (Gorishniy et al., 2024), we rank method A below method B if $|\text{B}_{\_}{\text{mean}}-\text{A}_{\_}{\text{mean}}|<\text{B}_{\_}{\text{stddev}}$ and B score is better. To further demonstrate the statistical significance of our findings, we ran a Tamhane’s T2 test of statistical significance for multiple comparisons. Results for Tamhane’s test are in Table 5
Table 5: Model ranks computed with Tamhane’s T2 statistical test. Ranking in not significantly altered compared to our simple testing procedure.
Methods Classification (ROC AUC ↑) Regression (RMSE ↓)
Clasical ML Baselines
XGBoost 3 3 1 1 3 2 1 2
LightGBM 3 3 2 2 4 2 2 1
CatBoost 2 4 3 2 3 1 3 2
RandomForest 4 3 7 4 6 7 8 7
Linear Model 9 3 8 3 7 6 9 8